Smith Corp. is in the process of liquidating and going out of business. The firm has $20,000 in cash, inventory totaling $1,000,000, accounts receivable of $200,000, plant and equipment with a $400,000 book value, and total liabilities of $700,000. It is estimated that the inventory can be disposed of in a liquidation sale for 65% of its cost, all but 25% of the accounts receivable can be collected, and plant and equipment can be sold for 30% of net book value.

1 ) Calculate the amount of cash that would be available to the stockholders if the accounts receivable are collected, the other assets are sold as described, and the liabilities are paid in full.
Answer #1

Ans:

Computation of cash available to stockholders:

Particular Amount $ Workings
Cash 20,000
Inventory sold 650,000 (1,000,000*65% = 650,000)
Accounts receivable collection 50,000 (200,000*25% = 50,000)
Plant and equipment sold 120,000 (400,000*30% = 120,000)
Total cash available before liabilities settled 840,000
Less: liabilities paid in full (700,000)
Cash available to stockholders 140,000

Total cash available to stockholders = $140,000.

If any doubts or queries please comment and clarify I'll explain ASAP.